# Discrete dark matter with light Dirac neutrinos

### v2

@@ VI Other detection prospects
In addition to the detection possibilities of DM and $\Delta N_{\rm eff}$, the model can have other detection aspects too. We briefly comment on such possibilities here.
@@
Due to the presence of several new scalars, admixtures of singlet and doublet, we can have different collider prospects as well. If scalar DM mass $(M_{\chi})$ is lighter than the SM Higgs mass $(M_{h})$ such that $h\rightarrow\chi\chi$ decay is allowed, it can contribute to invisible Higgs decay. Such invisible Higgs decay is currently constrained at Br$(h\rightarrow{\rm invisible})<0.14$ ATLAS:2022yvh . Production of $\mathcal{Z}_{2}$-odd scalars in colliders like the large hadron collider (LHC) can give rise to pure leptonic final states plus missing transverse energy (MET) Gustafsson:2012aj ; Datta:2016nfz , hadronic final states plus MET Poulose:2016lvz or a mixture of both. Such MET can arise from light Dirac neutrinos or scalar/fermion DM in our scenario. It is also possible to have more exotic signatures like tri-lepton plus MET Hashemi:2016wup and mono-jet Belyaev:2016lok ; Belyaev:2018ext . It should be noted that we also have new scalars in the model which are $\mathcal{Z}_{2}$-even. They can lead to additional final states, similar to generic two Higgs doublet models. The new charged scalars can also modify the diphoton decay width of the SM Higgs, currently constrained at $\frac{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm expt}}{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm SM}}=1.12\pm 0.09$ CMS:2021kom .
Another interesting detection prospect arises in the gravitational wave (GW) frontier. Spontaneous breaking of discrete symmetries like $A_{4}\times Z_{2}$, as we have in our model, can lead to formation of topological defects like domain walls (DW). While DW, if stable or allowed to dominate the energy density of the Universe, will be in conflict with cosmological observations, they can be made to disappear by introducing small bias, in the form of explicit symmetry breaking terms. Such annihilating or collapsing DW then generates stochastic GW which can be probed at ongoing or future experiments. GW signatures of minimal Dirac seesaw model was studied in earlier works Barman:2022yos ; Barman:2023fad while similar studies in the context of flavour symmetric models like $A_{4}$ can be found in Gelmini:2020bqg . The bias or explicit symmetry breaking terms introduced for DW disappearance may also lead to interesting indirect detection signatures of DM due to its decay into different standard model particles [^2] See King:2023ayw ; Borah:2024kfn for related discussions.. While the model has several detection prospects, it can also be falsified by future observation of neutrinoless double beta decay.
@@ VII Conclusion
The remnant $\mathcal{Z}_{2}$ symmetry allows either fermion or scalar DM candidates. While scalar DM is dominantly produced thermally due to gauge and scalar portal interactions, fermion DM production relies primarily on Yukawa portal interactions. The same Yukawa portal interactions can also lead to thermal or non-thermal production of right handed neutrinos, which contribute to the effective relativistic degrees of freedom $\Delta N_{\rm eff}$. To illustrate this, we consider three different benchmark scenarios of such Yukawa couplings and perform a detailed study of neutrino mass, mixing together with DM and $\Delta N_{\rm eff}$. We also check the scope of other detection prospects like LFV decays. Table 2 summarises our results. While one of the sub-classes of Yukawa couplings, namely Case-A, allows fermion singlet DM with observable LFV rates, the corresponding $\Delta N_{\rm eff}$ remains negligible. The other two sub-classes allow scalar DM only but with observable $\Delta N_{\rm eff}$ at future CMB experiments. The LFV rates however remain tiny in these two sub-classes. Due to the rich particle spectrum in terms of several scalar doublets and singlets, the model can also have rich collider phenomenology. On the other hand, spontaneous breaking of discrete symmetries can have interesting consequences in cosmology due to the formation of domain walls. Such walls, which disappear due to soft-breaking terms, can emit stochastic gravitational waves which can be probed at different ongoing and future experiments. The same soft-breaking terms can also lead to decay of DM, keeping indirect detection prospects alive. We leave a comprehensive analysis of such additional detection prospects of our model to future works.
@@ Appendix A $A_{4}$ Multiplication rules
All finite groups are completely characterized by means of a set of elements called generators of the group and a set of relations, so that all the elements of the group are given as product of the generators. The group $A_{4}$ consists of the even permutations of four objects and then contains 4!/2 = 12 elements. The generators are S and T with the relations $S^{2}=T^{3}=(ST)^{3}=\mathcal{I}$, then the elements are $1,S,T,ST,TS,T^{2},ST^{2},STS,TST,T^{2}S,TST^{2},T^{2}ST$.
@@
$1\displaystyle 1\quad S=1\quad T=$ 1
$1^{\prime}\quad S=1\quad T=\omega$
$1^{\prime\prime}\quad S=1\quad T=\omega^{2}$
$\omega=e^{i\frac{2\pi}{3}}$ is cube root of unity. In the basis, where $S$ is a real diagonal,
$S=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{pmatrix}\quad T=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$ (18)
@@
where ${\bf a}$ and ${\bf s}$ in the subscript corresponds to anti-symmetric and symmetric parts respectively. Now, if we have two triplets as $A=(a_{1},a_{2},a_{3})^{T}$ and $B=(b_{1},b_{2},b_{3})^{T}$ respectively, their direct product can be decomposed into the direct sum mentioned above. The product rule for this two triplets in the $S$ diagonal basis [^3] Here $S$ is a $3\times 3$ diagonal generator of $A_{4}$. can be written as He:2006dk ; Memenga:2013vc ; Ishimori:2010au
@@ Appendix B The scalar potential
The tree level scalar potential of our model can be written as
@@
$+\lambda_{\eta\chi 5}\big{[}(\eta^{\dagger}\eta)_{3_{2}}(\chi\chi)_{3_{1}}+{\rm h.c.}\big{]}+\lambda_{\eta H\chi}\big{[}(\chi\chi)_{3_{1}}(H^{\dagger}\eta)+{\rm h.c.}\big{]}$
@@
$\mathcal{M}^{2}_{S+}=\begin{pmatrix}M^{2}_{hh}&M^{2}_{hh_{1}}&M^{2}_{hx_{1}}\\ M^{2}_{hh_{1}}&M^{2}_{h_{1}h_{1}}&M^{2}_{h_{1}x_{1}}\\ M^{2}_{hx_{1}}&M^{2}_{h_{1}x_{1}}&M^{2}_{x_{1}x_{1}}\\ \end{pmatrix}_{3\times 3}$ (30)
@@
$M^{2}_{hh}=2\lambda_{H}v^{2},\,\,M^{2}_{hh_{1}}=(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})v_{\eta}v,\,\,M^{2}_{hx_{1}}=\lambda_{\chi H1}v_{\chi}v,$
$M^{2}_{h_{1}h_{1}}=2(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta},\,\,M^{2}_{h_{1}x_{1}}=(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))v_{\chi}v_{\eta},\,\,M^{2}_{x_{1}x_{1}}=2(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}.$
@@
$M^{2}_{h_{2}h_{2}}=$ $M_{h_{3}h_{3}}^{2}=\frac{1}{2}((-3\lambda_{\eta 2}+4\lambda_{\eta 3})v^{2}_{\eta}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}),$
$M^{2}_{x_{2}x_{2}}=$ $M_{x_{3}x_{3}}^{2}=\frac{1}{2}[-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\eta}^{2}+(-3\lambda_{\chi 2}+4\lambda_{\chi 3})v_{\chi}^{2}],$
$M^{2}_{h_{2}h_{3}}=$ $3\lambda_{\eta H4}vv_{\eta}$
$M^{2}_{h_{2}x_{2}}=$ $M_{h_{3}x_{3}}^{2}=\lambda_{\eta H\chi}vv_{\chi}+\lambda_{\eta\chi 4}v_{\eta}v_{\chi},$
$M_{h_{2}x_{3}}^{2}=$ $M_{h_{3}x_{2}}^{2}=\lambda_{\eta H\chi}vv_{\chi}$
$M_{x_{2}x_{3}}^{2}=$ $\lambda_{\eta H\chi}vv_{\eta}$
This mass squared matrix can similarly be diagonalised by a $4\times 4$ orthogonal matrix involving six angles, in general, leading to physical states $\zeta_{i}$. In the scalar DM case, the lightest eigenstate of the $4\times 4$ mass matrix given above becomes the DM candidate. Similar mass matrices can be written for both $\mathcal{Z}_{2}$-odd and $\mathcal{Z}_{2}$-even pseudoscalars as well.
$\mathcal{M}_{PS}^{2}=\begin{pmatrix}M^{2}_{AA}&M^{2}_{AA_{1}}&0&0\\ M^{2}_{AA_{1}}&M^{2}_{A_{1}A_{1}}&0&0\\ 0&0&M^{2}_{A_{2}A_{2}}&M^{2}_{A_{2}A_{3}}\\ 0&0&M^{2}_{A_{2}A_{3}}&M^{2}_{A_{3}A_{3}}\end{pmatrix}_{4\times 4}$ (32)
where
$M^{2}_{AA}=$ $-2\lambda_{\eta H3}v_{\eta}^{2},\quad M^{2}_{A_{1}A_{1}}=-2\lambda_{\eta H3}v^{2},\quad M^{2}_{AA_{1}}=2\lambda_{\eta H3}vv_{\eta},$
$M_{A_{2}A_{2}}^{2}=$ $M_{A_{3}A_{3}}^{2}=\frac{1}{2}[-4\lambda_{\eta H3}v^{2}-(3\lambda_{\eta 2}+8\lambda_{\eta 5})v_{\eta}^{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}],$
@@
Similarly for the charged scalars, the mass matrix squared in the ($h^{\pm},\eta_{1}^{\pm},\eta_{2}^{\pm},\eta_{3}^{\pm}$) basis can be expressed as
$\mathcal{M}^{2}_{CS}=\begin{pmatrix}M^{2}_{h^{\pm}h^{\mp}}&M^{2}_{h^{\pm}\eta_{1}^{\mp}}&0&0\\ M^{2}_{h^{\pm}\eta_{1}^{\mp}}&M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}&0&0\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{2}^{\mp}}&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}&M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}\end{pmatrix}_{4\times 4}$ (33)
$M^{2}_{h^{\pm}h^{\mp}}=$ $-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v_{\eta}^{2}}{2},\quad M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2},$
$M_{\eta_{2}^{\pm}\eta_{2}^{\mp}}^{2}=$ $M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2}-3\lambda_{\eta 2}\frac{v_{\eta}^{2}}{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})\frac{v_{\chi}^{2}}{2},$
$M^{2}_{h^{\pm}\eta_{1}^{\mp}}=$ $(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{vv_{\eta}}{2},\quad M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}=\lambda_{\eta H4}vv_{\eta}.$

### v3

@@ VI Other detection prospects
In addition to the detection possibilities of DM and $\Delta{\rm N}_{\rm eff}$, the model can have other detection aspects too. We briefly comment on such possibilities in this section.
@@
Due to the presence of several new scalars, admixtures of singlet and doublet, we can have different collider prospects as well. If scalar DM mass $(M_{\chi})$ is lighter than the SM Higgs mass $(M_{h})$ such that $h\rightarrow\chi\chi$ decay is allowed, it can contribute to invisible Higgs decay. Such invisible Higgs decay is currently constrained at Br$(h\rightarrow{\rm invisible})<0.14$ ATLAS:2022yvh . Production of $\mathcal{Z}_{2}$-odd scalars in colliders like the large hadron collider (LHC) can give rise to pure leptonic final states plus missing transverse energy (MET) Gustafsson:2012aj ; Datta:2016nfz , hadronic final states plus MET Poulose:2016lvz or a mixture of both. Such MET can arise from light Dirac neutrinos or scalar/fermion DM in our scenario. It is also possible to have more exotic signatures like tri-lepton plus MET Hashemi:2016wup and mono-jet Belyaev:2016lok ; Belyaev:2018ext . It should be noted that we also have new scalars in the model which are $\mathcal{Z}_{2}$-even. They can lead to additional final states, similar to generic two-Higgs doublet models. The new charged scalars can also modify the diphoton decay width of the SM Higgs, currently constrained at $\frac{{\rm Br}(h\rightarrow\gamma\gamma)_{\rm expt}}{{\rm Br}(h\rightarrow\gamma\gamma)_{\rm SM}}=1.12\pm 0.09$ CMS:2021kom .
Another interesting detection prospect arises in the gravitational wave (GW) frontier. Spontaneous breaking of discrete symmetries like $A_{4}\times Z_{2}$, as we have in our model, can lead to the formation of topological defects like domain walls (DW). While DW, if stable or allowed to dominate the energy density of the Universe, will be in conflict with cosmological observations, they can be made to disappear by introducing small bias, in the form of explicit symmetry breaking terms. Such annihilating or collapsing DW then generates stochastic GW which can be probed at ongoing or future experiments. GW signatures of minimal Dirac seesaw model was studied in earlier works Barman:2022yos ; Barman:2023fad while similar studies in the context of flavour symmetric models like $A_{4}$ can be found in Gelmini:2020bqg . The bias or explicit symmetry breaking terms introduced for DW disappearance may also lead to interesting indirect detection signatures of DM due to its decay into different standard model particles [^2] See King:2023ayw ; Borah:2024kfn for related discussions.. While the model has several detection prospects, it can also be falsified by future observation of neutrinoless double beta decay.
@@ VII Conclusion
The remnant $\mathcal{Z}_{2}$ symmetry allows either fermion or scalar DM candidates. While scalar DM is dominantly produced thermally due to gauge and scalar portal interactions, fermion DM production relies primarily on Yukawa portal interactions. The same Yukawa portal interactions can also lead to thermal or non-thermal production of right handed neutrinos, which contribute to the effective relativistic degrees of freedom $\Delta{\rm N}_{\rm eff}$. To illustrate this, we consider three different benchmark scenarios of such Yukawa couplings and perform a detailed study of neutrino mass, mixing together with DM and $\Delta{\rm N}_{\rm eff}$. We also check the scope of other detection prospects like LFV decays. Table 2 summarises our results. While one of the sub-classes of Yukawa couplings, namely Case-A, allows fermion singlet DM with observable LFV rates, the corresponding $\Delta{\rm N}_{\rm eff}$ remains negligible. The other two sub-classes allow scalar DM only but with observable $\Delta{\rm N}_{\rm eff}$ at future CMB experiments. The LFV rates, however, remain tiny in these two sub-classes. Due to the rich particle spectrum in terms of several scalar doublets and singlets, the model can also have rich collider phenomenology. On the other hand, spontaneous breaking of discrete symmetries can have interesting consequences in cosmology due to the formation of domain walls. Such walls, which disappear due to soft-breaking terms, can emit stochastic gravitational waves, which can be probed in different ongoing and future experiments. The same soft-breaking terms can also lead to the decay of DM, keeping indirect detection prospects alive. We leave a comprehensive analysis of such additional detection prospects of our model to future works.
@@ Appendix A $A_{4}$ Multiplication rules
All finite groups are completely characterized by a set of elements called generators of the group and a set of relations, so that all the elements of the group are given as products of the generators. The group $A_{4}$ consists of the even permutations of four objects and then contains 4!/2 = 12 elements. The generators are S and T with the relations $S^{2}=T^{3}=(ST)^{3}=\mathcal{I}$, then the elements are $1,S,T,ST,TS,T^{2},ST^{2},STS,TST,T^{2}S,TST^{2},T^{2}ST$.
@@
$1\quad S=1\quad T=1;$
$1^{\prime}\quad S=1\quad T=\omega;$
$1^{\prime\prime}\quad S=1\quad T=\omega^{2};$
$\omega=e^{i\frac{2\pi}{3}}$, is cube root of unity. In the basis, where $S$ is a real diagonal,
$S=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{pmatrix}_{3\times 3}\quad T=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix}_{3\times 3}$ (18)
@@
where ${\bf 1}$ and ${\bf 2}$ in the subscript correspond to anti-symmetric and symmetric parts, respectively. Now, if we have two triplets as $A=(a_{1},a_{2},a_{3})^{T}$ and $B=(b_{1},b_{2},b_{3})^{T}$ respectively, their direct product can be decomposed into the direct sum mentioned above. The product rule for this two triplets in the $S$ diagonal basis [^3] Here $S$ is a $3\times 3$ diagonal generator of $A_{4}$. can be written as He:2006dk ; Memenga:2013vc ; Ishimori:2010au
@@ Appendix B The scalar potential
The tree-level scalar potential of our model can be written as
@@
$+\lambda_{\eta\chi 5}\big{[}(\eta^{\dagger}\eta)_{3_{2}}(\chi\chi)_{3_{1}}+{\rm h.c.}\big{]}+\lambda_{\eta H\chi}\big{[}(\chi\chi)_{3_{1}}(H^{\dagger}\eta)+{\rm h.c.}\big{]},$
@@
$\mathcal{M}^{2}_{S+}=\begin{pmatrix}M^{2}_{hh}&M^{2}_{hh_{1}}&M^{2}_{hx_{1}}\\ M^{2}_{hh_{1}}&M^{2}_{h_{1}h_{1}}&M^{2}_{h_{1}x_{1}}\\ M^{2}_{hx_{1}}&M^{2}_{h_{1}x_{1}}&M^{2}_{x_{1}x_{1}}\\ \end{pmatrix}_{3\times 3},$ (30)
@@
$M^{2}_{hh}=2\lambda_{H}v^{2};\,\,M^{2}_{hh_{1}}=(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})v_{\eta}v;\,\,M^{2}_{hx_{1}}=\lambda_{\chi H1}v_{\chi}v;$
$M^{2}_{h_{1}h_{1}}=2(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta};\,\,M^{2}_{h_{1}x_{1}}=(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))v_{\chi}v_{\eta};\,\,M^{2}_{x_{1}x_{1}}=2(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}.$
@@
where,
$M^{2}_{h_{2}h_{2}}=$ $M_{h_{3}h_{3}}^{2}=\frac{1}{2}((-3\lambda_{\eta 2}+4\lambda_{\eta 3})v^{2}_{\eta}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2});$
$M^{2}_{x_{2}x_{2}}=$ $M_{x_{3}x_{3}}^{2}=\frac{1}{2}[-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\eta}^{2}+(-3\lambda_{\chi 2}+4\lambda_{\chi 3})v_{\chi}^{2}];$
$M^{2}_{h_{2}h_{3}}=$ $3\lambda_{\eta H4}vv_{\eta};$
$M^{2}_{h_{2}x_{2}}=$ $M_{h_{3}x_{3}}^{2}=\lambda_{\eta H\chi}vv_{\chi}+\lambda_{\eta\chi 4}v_{\eta}v_{\chi};$
$M_{h_{2}x_{3}}^{2}=$ $M_{h_{3}x_{2}}^{2}=\lambda_{\eta H\chi}vv_{\chi};$
$M_{x_{2}x_{3}}^{2}=$ $\lambda_{\eta H\chi}vv_{\eta}.$
This mass squared matrix can similarly be diagonalised by a $4\times 4$ orthogonal matrix involving six angles, in general, leading to physical states. In the scalar DM case, the lightest eigenstate of the $4\times 4$ mass matrix given above becomes the DM candidate. Similar mass matrices can be written for both $\mathcal{Z}_{2}$-odd and $\mathcal{Z}_{2}$-even pseudoscalars as well.
$\mathcal{M}_{PS}^{2}=\begin{pmatrix}M^{2}_{AA}&M^{2}_{AA_{1}}&0&0\\ M^{2}_{AA_{1}}&M^{2}_{A_{1}A_{1}}&0&0\\ 0&0&M^{2}_{A_{2}A_{2}}&M^{2}_{A_{2}A_{3}}\\ 0&0&M^{2}_{A_{2}A_{3}}&M^{2}_{A_{3}A_{3}}\end{pmatrix}_{4\times 4},$ (32)
@@
$M^{2}_{AA}=$ $-2\lambda_{\eta H3}v_{\eta}^{2};\quad M^{2}_{A_{1}A_{1}}=-2\lambda_{\eta H3}v^{2};\quad M^{2}_{AA_{1}}=2\lambda_{\eta H3}vv_{\eta};$
$M_{A_{2}A_{2}}^{2}=$ $M_{A_{3}A_{3}}^{2}=\frac{1}{2}[-4\lambda_{\eta H3}v^{2}-(3\lambda_{\eta 2}+8\lambda_{\eta 5})v_{\eta}^{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}];$
@@
Similarly, for the charged scalars, the mass matrix squared in the ($h^{\pm},\eta_{1}^{\pm},\eta_{2}^{\pm},\eta_{3}^{\pm}$) basis can be expressed as
$\mathcal{M}^{2}_{CS}=\begin{pmatrix}M^{2}_{h^{\pm}h^{\mp}}&M^{2}_{h^{\pm}\eta_{1}^{\mp}}&0&0\\ M^{2}_{h^{\pm}\eta_{1}^{\mp}}&M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}&0&0\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{2}^{\mp}}&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}&M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}\end{pmatrix}_{4\times 4},$ (33)
$M^{2}_{h^{\pm}h^{\mp}}=$ $-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v_{\eta}^{2}}{2};\quad M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2};$
$M_{\eta_{2}^{\pm}\eta_{2}^{\mp}}^{2}=$ $M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2}-3\lambda_{\eta 2}\frac{v_{\eta}^{2}}{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})\frac{v_{\chi}^{2}}{2};$
$M^{2}_{h^{\pm}\eta_{1}^{\mp}}=$ $(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{vv_{\eta}}{2};\quad M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}=\lambda_{\eta H4}vv_{\eta}.$
@@
For illustrative purposes, we use the following benchmark parameters related to the scalar sector to estimate the scalar masses numerically while being in agreement with the loop functions used in neutrino mass analysis: $v=174$ GeV, $v_{\eta}=174$ GeV, $v_{\chi}=4800$ GeV, $\lambda_{H}=0.188$, $\lambda_{\eta 1}=4.9$, $\lambda_{\eta_{2}}=4.9$, $\lambda_{\eta 3}=0.4$, $\lambda_{\eta 5}=0.6$, $\lambda_{\eta H1}=5.3$, $\lambda_{\eta H2}=-0.8$, $\lambda_{\eta H3}=-3$, $\lambda_{\eta H4}=1$, $\lambda_{\eta\chi 1}=0.001$, $\lambda_{\eta\chi 2}=-0.5$, $\lambda_{\eta\chi 3}=0.48$, $\lambda_{\eta\chi 4}=0.1$, $\lambda_{\chi 1}=0.1$, $\lambda_{\chi 2}=0.5$, $\lambda_{\chi 3}=0.8$, $\lambda_{\chi H1}=0.02$, $\lambda_{\eta H\chi}=0.5$. The physical scalar masses for the chosen benchmark are shown in table 3.
Scalars Masses Loop functions
$Z_{2}$-even neutral scalars $M_{1+}=125$ GeV $M_{2+}=775$ GeV $M_{3+}=5258$ GeV $\mathcal{F}_{1}=1\times 10^{-3}\newline \mathcal{F}_{2}=6.1\times 10^{-4}$
$\mathcal{F}_{3}=5.96\times 10^{-4}$
$Z_{2}$-odd neutral scalars $M_{1-}=603$ GeV $M_{2-}=4423$ GeV $M_{3-}=733$ GeV $M_{4-}=4437$ GeV
$Z_{2}$-odd charged scalars $M_{1c}=785$ GeV $M_{2c}=1169$ GeV $M_{3c}=1154$ GeV
$Z_{2}$-odd pseudo-scalars $M_{1ps}=738$ GeV $M_{2ps}=866$ GeV $M_{3ps}=876$ GeV
Table 3: Numerical estimation for the scalar sector with associated loop functions. We use the notation $M_{i}$ for the respective mass eigenvalues of scalar mass matrices shown in Eq. (30), (31), (32), (33).